The average price of a particular item was $4.30 in 1996.  The average price in 2017 was $8.97.  The CPI in 1996 was 156.9.  The CPI in 2017 was 237.0.  After adjusting for inflation, has the real price of a movie ticket increased? (show work)
Answer #1

answer

Yes, it is increased.

-----

real Price in year x =( nominal price/CPI)*100

real price in 1996=(4.30/156.9)*100=2.74059911=2.74

real price in 2017=(8.97/237)*100=3.78481013=3.78

the real price is increased from 2.74 to 3.78 from 1996 to 2017.